3. Consider a normal distribution with a mean of 500 and a standard deviation of 50.
a. Below what value can we expect to have the lowest 20%?
b. Between what values can we expect to find the middle 80%?
1
Expert's answer
2020-10-22T17:49:37-0400

"\mu =500, \sigma=50"


a) "P(X<N)=20\%=0.2"

Z score is -0.84 then

"\frac{N-500}{50}=-0.84 \implies N=500-50*0.84=458"


b) "P(N_1<X<N_2)=80\%"

"P(X<N_2)=90\%=0.9\implies" Z score is 1.28

"\frac{N_2-500}{50}=1.28 \implies N_2=500+50*1.28=564"

"P(X<N_1)=10\%=0.1\implies" Z score is -1.28

"\frac{N_1-500}{50}=-1.28 \implies N_1=500-50*1.28=436"


Answer:

a) 458

b) between 436 and 564